Greenwood Company manufactures two products—13,000 units of Product Y and 5,000 units of Product Z. The company uses a plantwide overhead rate based on direct labor-hours. It is considering implementing an activity-based costing (ABC) system that allocates all of its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products Y and Z: Activity Cost Pool Activity Measure Estimated Overhead Cost Expected Activity Machining Machine-hours $ 242,400 12,000 MHs Machine setups Number of setups $ 114,400 220 setups Production design Number of products $ 86,000 2 products General factory Direct labor-hours $ 302,400 12,000 DLHs Activity Measure Product Y Product Z Machining 8,200 3,800 Number of setups 40 180 Number of products 1 1 Direct labor-hours 8,200 3,800

Using the plantwide overhead rate, how much manufacturing overhead cost is allocated to Product Y and Product Z? What is the activity rate for the Machining activity cost pool? What is the activity rate for the Machine Setups activity cost pool? What is the activity rate for the Product Design activity cost pool? What is the activity rate for the General Factory activity cost pool? Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Y? Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Z? Using the plantwide overhead rate, what percentage of the total overhead cost is allocated to Product Y and Product Z? Using the ABC system, what percentage of the Machining costs is assigned to Product Y and Product Z? Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z? Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z? Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z?
Answer #1

Estimated overhead cost= $242400+114400+86000+302400= $745200

Total direct labor hours= 8200+3800= 12000 labor hours

Plantwide overhead rate= Estimated overhead cost/Total direct labor hours

= $745200/12000= $62.10 per labor hour

1) Using the plantwide overhead rate, how much manufacturing overhead cost is allocated to Product Y and Product Z?

Product Y= 8200*$62.10= $509220

Product Z= 3800*$62.10= $235980

2)  What is the activity rate for the Machining activity cost pool?

Activity rate= Estimated overhead cost/Estimated activity

Machining cost pool= $242400/12000= $20.20 per machine hour

3) What is the activity rate for the Machine Setups activity cost pool?

Machine setups activity cost pool= $114400/220= $520 per machine setups

4) What is the activity rate for the Product Design activity cost pool?

Product Design= $86000/2= $43000 per product

5) What is the activity rate for the General Factory activity cost pool?

General factory= $302400/12000= $25.20 per direct labor hour

6) Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Y?

Product Y
Machining (8200*$20.20) $165640
Machine setup (40*$520) 20800
Product design (1*$43000) 43000
General factory ($25.20*8200) 206640
Total manufacturing overhead cost $436080

7) Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Z?

Product Y
Machining (3800*$20.20) $76760
Machine setup (180*$520) 93600
Product design (1*$43000) 43000
General factory ($25.20*3800) 95760
Total manufacturing overhead cost $309120

8) Using the plantwide overhead rate, what percentage of the total overhead cost is allocated to Product Y and Product Z?

Total manufacturing overhead cost= $242400+114400+86000+302400= $745200

Total overhead cost allocated to Product Y= $509220*100/745200= 68.33%

Total overhead cost allocated to Product Y= $235980*100/745200= 31.67%

9) Using the ABC system, what percentage of the Machining costs is assigned to Product Y and Product Z?

Machining costs assigned to Product Y= $165640*100/242400= 68.33%

Machining costs assigned to Product Z= $76760*100/242400= 31.67%

10) Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z?

Machine Setups cost assigned to Product Y= $20800*100/114400= 18.18%

Machine Setups cost assigned to Product Z= $93600*100/114400= 81.82%

11) Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z?

Product Design cost assigned to Product Y= $43000*100/86000= 50%

Product Design cost assigned to Product Z= $43000*100/86000= 50%

12) Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z?

General Factory cost assigned to Product Y= $206640*100/302400= 68.33%

General Factory cost assigned to Product Z= $95760*100/302400= 31.67%